OpenStudy (anonymous):

how to find the gradient of a line when given y+'half' (1/2)=1

OpenStudy (anonymous):

gradient of y+1/2=1 ??

OpenStudy (anonymous):

if y+1/2=1 is parrellel to y=2x+1 ive already worked out y=2x+1 just dont know how to do the second part of the question which is y+1/2=1 key- 1/2 =half